# Let’s Be Self-generated via Step by Step: A Curriculum Learning Approach to Automated Reasoning with Large Language Models

### v1

@@ Abstract
While Chain of Thought (CoT) prompting approaches have significantly consolidated the reasoning capabilities of large language models (LLMs), they still face limitations that require extensive human effort or have performance needs to be improved. Existing endeavors have focused on bridging these gaps; however, these approaches either hinge on external data and cannot completely eliminate manual effort, or they fall short in effectively directing LLMs to generate high-quality exemplary prompts. To address the said pitfalls, we propose a novel prompt approach for automatic reasoning named LBS3, inspired by curriculum learning which better reflects human learning habits. Specifically, LBS3 initially steers LLMs to recall easy-to-hard proxy queries that are pertinent to the target query. Following this, it invokes a progressive strategy that utilizes exemplary prompts stemmed from easy-proxy queries to direct LLMs in solving hard-proxy queries, enabling the high-quality of the proxy solutions. Finally, our extensive experiments in various reasoning-intensive tasks with varying open- and closed-source LLMs show that LBS3 achieves strongly competitive performance compared to the SOTA baselines. Our code is here: https://anonymous.4open.science/r/LBS3-B926.
@@ 1 Introduction
With super-sized training corpora and computational cluster resources, Large Language Models (LLMs) have demonstrated emergent capabilities, thus enabling state-of-the-art performance in a wide range of natural language tasks (Wei et al., 2022a; Brown et al., 2020; Chowdhery et al., 2023; Liang et al., 2022; Qin et al., 2023; Wei et al., 2023; Touvron et al., 2023). However, directly applying LLMs to complex reasoning tasks (e.g., mathematical reasoning, commonsense reasoning, etc.) in a naive manner presents significant challenges (Yin et al., 2023b; Wei et al., 2022b; Kojima et al., 2022). For instance, the performance may be inadequate when simply feeding queries or using few-shot query-answer pairs in in-context learning (ICL) approaches for these kinds of tasks. Recent studies have shed light on that prompting LLMs to generate multiple reasoning steps (i.e., rationale) can markedly enhance their ability to reason, resulting in the development of the chain-of-thought (CoT) prompting (Wei et al., 2022b; Kojima et al., 2022; Zhou et al., 2022b; Wang et al., 2022b; Aggarwal et al., 2023a; Chen et al., 2024a; Yao et al., 2024; Zou et al., 2023; Yu et al., 2024; Besta et al., 2024). Current CoT prompting approaches fall broadly into two categories, namely Few-Shot CoT (FS-CoT) (Wei et al., 2022b) and Zero-Shot CoT(ZS-CoT) (Kojima et al., 2022). Among them, FS-CoT involves providing LLMs with few task-specific context exemplars of query-rationale-answer triplets tied to the target query to prompt the generation of reasoning steps; ZS-CoT instead stimulates LLMs’ reasoning capabilities by furnishing general trigger instructions (such as "Let’s think step by step") appended to the target query.
@@
Despite their considerable success, obstacles persist in the field of prompt engineering research that plague real-world applications. FS-CoT, for example, delivers well-crafted exemplary prompts but at the cost of labor-intensive manual annotations. To mitigate this, some efforts have been made to enhance the quality of exemplary prompts by retrieving the most relevant, complex and diverse existing queries or exemplars for the target task, which is achieved by tapping into external sources related to the task at hand, such as datasets or corpora, and employing various pre-defined similarity metrics (Liu et al., 2021; Rubin et al., 2021; Fu et al., 2022; Ye et al., 2022; Su et al., 2022; Wu et al., 2022; Ye & Durrett, 2023; Diao et al., 2023; Wan et al., 2023). Nevertheless, the required external sources these approaches rely on may not always be available in practice, and they may not completely obviate the need for manual labeling. Moreover, while ZS-CoT offers versatility, its performance often lags behind FS-CoT in a variety of complex reasoning tasks.
To overcome the aforementioned issues, recent initiatives (e.g., Self-ICL (Chen et al., 2023), Auto-ICL (Yang et al., 2023) and Analogical Prompting (Yasunaga et al., 2023)) work on bootstrapping LLMs to self-generate few-shot new proxy queries that are relevant and diverse to the target query, along with constructing corresponding exemplary prompts of triplets, thereby augmenting their capabilities to tackle reasoning tasks. Essentially, these methods draw parallels with the concept of analogical reasoning in psychology, where individuals rely on past related experiences to approach new problems Vosniadou & Ortony (1989). The underlying insight behind them is that pre-trained LLMs (such as Llama3, GPT-3.5 and GPT-4.0, etc.) have amassed a wealth of knowledge that equips them to fulfill various reasoning tasks. However, we observed that merely prompting LLMs to recall experiences related to the target queries may lead to the generation of proxy queries that are as difficult as the target queries themselves, along with corresponding incorrect proxy solutions, misguiding the resolution of the target queries, as exemplified in Fig. 1. See the related works in Appendix A for more details.
The issues mentioned above motivate us to propose a novel automatic reasoning prompt approach, coined LBS3, which is inspired by curriculum learning that mirrors the progressive nature of human learning styles. The idea of curriculum learning (Bengio et al., 2009; Cornacchia & Mossel, 2023) has been widely applied in the field of artificial intelligence, emulating the human learning process from easy to hard tasks (Campos, 2021; Maharana & Bansal, 2022; Huang et al., 2020; Kong et al., 2021; Zhu et al., 2022; Li et al., 2021; Soviany et al., 2022; Xu et al., 2020). Thus, LBS3 critically aims to 1) guide LLMs to generate easy- (or hard-) proxy queries related to the target query and 2) enhance the effectiveness of the solutions for these proxy queries to benefit that for the target query. For the former, diverging from existing approaches (e.g., Self-ICL and Auto-ICL) that generate proxy queries in one pass, we develop a two-stage framework of generation for proxy queries. Specifically, we first prompt LLMs with instructions like "Generate $n_{1}$ different new relevant problems that are easier to solve than the example problem below." to generate simpler proxy queries than the given query, that is, easy-proxy queries. Then, we instruct LLMs to formulate analogical proxy queries for the given query, which are the hard-proxy queries, by using instructions like "Generate $n_{2}$ different new problems that are analogous to the example problem below." Note that $n_{1}$ and $n_{2}$ denote the number of proxy queries generated. For the latter, we initially leverage LLMs to solve each easy-proxy query independently in the ZS-CoT manner, creating corresponding triplet exemplary prompt. Subsequently, we combine these prompts with each hard-proxy query and generate solutions in the FS-CoT fashion. Ultimately, we amalgamate all constructed exemplary prompts with the given query and derive the target solution in the FS-CoT manner. We modularly outline the generic framework of the reasoning process for LBS3 in Fig. 2.
@@
One of the advantages for the proposed approach is that LBS3 explicitly distinguishes between easy- and hard-proxy queries, and ensures that the difficulty of solving proxy queries does not exceed that of the given query. Additionally, in contrast to existing approaches that tackle each proxy query from scratch, we adopt a progressive strategy to harness exemplary prompts derived from easy-proxy queries to guide the generation of solution for hard ones, thereby alleviating the accumulation of errors arises from reasoning ab initio (Yu et al., 2024). To the best of our knowledge, our work is the pioneering attempt to emulate the idea of curriculum learning, aiming to investigate how LLMs can self-generate few-shot exemplary prompts to facilitate the reasoning process.
@@ 3.2 Progressive Strategy of Solving Queries
Algorithm 1 The pseudocode of LBS3 approach given one query $[q]$.
Input: $[q]$: the target query, $LLM$: large language model, $n_{1}$: the number of easy-proxy queries, $n_{2}$: the number of hard-proxy queries
1:Initial modules: SPG, APG and RAG
2:# Stage 1:
3:prompt_spg=SPG.format($n_{1}$, problem=$[q]$)
[…]
5:exem_sa = {}
6:for $[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]$ in $\{[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]\}_{i=1}^{n_{1}}$ do
7:     prompt_rag-z = RAG-Z.format(problem=$[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]$)
8:     $[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{i}}]$= $LLM($prompt_rag-z))))
9:     exem_sa = exem_sa $\bigcup$ $\{[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}][{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{i}}]\}$
10:end for
11:# Stage 2:
12:prompt_apg = APG.format($n_{2}$, problem=$[q]$)
[…]
18:end for
19:# Stage 3:
20:prompt_rag-f = RAG-F.format(exemplars=exem_sa, problem=$[q]$)
21:$[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a}]$ = $LLM($prompt_rag-f))))
Output: $[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a}]$
The primary advantage of the above strategy is its effectiveness in enhancing the solutions for hard-proxy queries. To be specific, the easy-to-solve $\{[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]\}_{i=1}^{n_{1}}$ ensures that the corresponding exemplary prompts may be correct with high confidence. Meanwhile, (Chen et al., 2023; Yang et al., 2023; Yasunaga et al., 2023) have shown that using solved proxy queries related or analogous to $[q]$ as exemplary prompts can effectively improve the solution for $[q]$. However, when the difficulty of solving query $[q]$ is high, the generated proxy queries are likely to have comparable challenging, resulting in low-quality exemplary prompts. Therefore, adopting our proposed progressive strategy can alleviate the cumulative errors associated with solving hard-proxy queries from scratch (Yu et al., 2024), thereby enhancing the quality of their solutions. In addition, we find that using already solved hard-proxy queries as additional exemplary prompts for solving the next hard-proxy query can further strengthen the solution to $[q]$, see Alg. 1 and Section 4.3.3 for more details and empirical justification.
4 Experiments
4.1 Experimental Settings
4.2 Results Comparison
Furthermore, the baselines Self-ICL and Auto-ICL uniformly beat ZS-CoT in terms of average accuracy and surpassed FS-CoT in most cases. This result highlights that guiding LLMs to autonomously generate proxy exemplars relevant to a given query can effectively improve their reasoning capabilities. Additionally, the baseline Ana-Pro consistently underperforms other competitors w.r.t. average accuracy, including ZS-CoT, on open-source LLMs, yet consistently outstrips ZS-CoT w.r.t. average accuracy on closed-source LLMs. The said result confirms the high requirements imposed by the Ana-Pro approach on LLMs for following instructions and generating responses. It is worth noting that the open-source model Llama3-70B-Instruct achieves the best accuracy across all prompting approaches and even shows significant performance gains compared with the closed-source model GPT-4.0-turbo in most cases. We speculate that this phenomenon occurs because we only considered a limited set of reasoning tasks, and thus closed-source models may show stronger generalization capabilities on more and wider tasks.
@@ 4.2 Results Comparison
LBS3 94.9 64.2 89.8 86.6 86.0 74.8 82.8 96.0 84.38
@@
4.3 Ablation Study
In this section, we carefully demonstrate the efficacy and indispensability of the core components in our approach on Qwen1.5-14B-Chat, Llama3-70B-Instruct, and GPT-3.5-turbo over diverse benchmarks. These components include the two-stage framework for self-generating proxy queries from easy to hard, as well as the progressive strategy employed to solve them.
4.3.1 Comparison of the number for easy and hard proxy exemplars
We look into the impacts of different hyperparameter combinations ($n_{1}$, $n_{2}$) within the two-stage framework for self-generating proxy queries of LBS3 across various benchmarks, including GSM8K, SQA, CQA, BBH-idfo, BBH-raco, and BBH-ts. For clarity, assume that the number of proxy exemplars $n$ is $3$, with both $n_{1}$ and $n_{2}$ taking values from $\{0,1,2,3\}$. Note that since $n=n_{1}+n_{2}$, we opt to only consider $n_{1}$, then $n_{2}=n-n_{1}$. When $n_{1}=1$ and $n_{2}=2$, this indicates that LBS3 is tasked with generating one easy-proxy exemplar and two hard-proxy ones. We report the performance of LBS3 corresponding to varying $n_{1}$ on different models and benchmarks in Fig. 3.
@@ 4.3.1 Comparison of the number for easy and hard proxy exemplars
4.3.2 LBS3 with Self-ICL and Auto-ICL
4.3.3 Utility of Progressive Strategy

### v3

@@ Abstract
While Chain of Thought (CoT) prompting approaches have significantly consolidated the reasoning capabilities of large language models (LLMs), they still face limitations that require extensive human effort or have performance needs to be improved. Existing endeavors have focused on bridging these gaps; however, these approaches either hinge on external data and cannot completely eliminate manual effort, or they fall short in effectively directing LLMs to generate high-quality exemplary prompts. To address the said pitfalls, we propose a novel prompt approach for automatic reasoning named LBS3, inspired by curriculum learning which better reflects human learning habits. Specifically, LBS3 initially steers LLMs to recall easy-to-hard proxy queries that are pertinent to the target query. Following this, it invokes a progressive strategy that utilizes exemplary prompts stemmed from easy-proxy queries to direct LLMs in solving hard-proxy queries, enabling the high-quality of the proxy solutions. Finally, our extensive experiments in various reasoning-intensive tasks with varying open- and closed-source LLMs show that LBS3 achieves strongly competitive performance compared to the SOTA baselines. Our code is here: https://anonymous.4open.science/r/DFRD-0C83/.
Let’s Be Self-generated via Step by Step: A Curriculum Learning Approach to Automated Reasoning with Large Language Models
Kangyang Luo♠♢, Zichen Ding, Zhenmin Weng, Lingfeng Qiao, Meng Zhao Xiang Li, Di Yin, Jinlong Shu East China Normal University   Youtu Lab, Tencent   Shanghai Normal University
@@ 1 Introduction
With super-sized training corpora and computational cluster resources, Large Language Models (LLMs) have demonstrated emergent capabilities, thus enabling state-of-the-art performance in a wide range of natural language tasks (Wei et al., 2022a; Brown et al., 2020; Chowdhery et al., 2023; Liang et al., 2022; Qin et al., 2023; Wei et al., 2023; Touvron et al., 2023). However, directly applying LLMs to complex reasoning tasks (e.g., mathematical reasoning, commonsense reasoning, etc.) in a naive manner presents significant challenges (Yin et al., 2023b; Wei et al., 2022b; Kojima et al., 2022). For instance, the performance may be inadequate when simply feeding queries or using few-shot query-answer pairs in in-context learning (ICL) approaches for these kinds of tasks. Recent studies have shed light on that prompting LLMs to generate multiple reasoning steps (i.e., rationale) can markedly enhance their ability to reason, resulting in the development of the chain-of-thought (CoT) prompting (Wei et al., 2022b; Kojima et al., 2022; Zhou et al., 2022b; Wang et al., 2022b; Aggarwal et al., 2023a; Chen et al., 2024a; Yao et al., 2024; Zou et al., 2023; Yu et al., 2024b; Besta et al., 2024). Current CoT prompting approaches fall broadly into two categories, namely Few-Shot CoT (FS-CoT) (Wei et al., 2022b) and Zero-Shot CoT(ZS-CoT) (Kojima et al., 2022). Among them, FS-CoT involves providing LLMs with few task-specific context exemplars of query-rationale-answer triplets tied to the target query to prompt the generation of reasoning steps; ZS-CoT instead stimulates LLMs’ reasoning capabilities by furnishing general trigger instructions (such as "Let’s think step by step") appended to the target query.
@@
Despite their considerable success, obstacles persist in the field of prompt engineering research that plague real-world applications. FS-CoT, for example, delivers well-crafted exemplary prompts but at the cost of labor-intensive manual annotations. To mitigate this, some efforts have been made to enhance the quality of exemplary prompts by retrieving the most relevant, complex and diverse existing queries or exemplars for the target task, which is achieved by tapping into external sources related to the task at hand, such as datasets or corpora, and employing various pre-defined similarity metrics (Liu et al., 2021; Rubin et al., 2021; Fu et al., 2022; Ye et al., 2022; Su et al., 2022; Wu et al., 2022; Ye and Durrett, 2023; Diao et al., 2023; Wan et al., 2023). Nevertheless, the required external sources these approaches rely on may not always be available in practice, and they may not completely obviate the need for manual labeling. Moreover, while ZS-CoT offers versatility, its performance often lags behind FS-CoT in a variety of complex reasoning tasks.
@@
To overcome the aforementioned issues, recent initiatives (e.g., Self-ICL (Chen et al., 2023), Auto-ICL (Yang et al., 2023) and Analogical Prompting (Yasunaga et al., 2023)) work on bootstrapping LLMs to self-generate few-shot new proxy queries that are relevant and diverse to the target query, along with constructing corresponding exemplary prompts of triplets, thereby augmenting their capabilities to tackle reasoning tasks. Essentially, these methods draw parallels with the concept of analogical reasoning in psychology, where individuals rely on past related experiences to approach new problems (Vosniadou and Ortony, 1989). The underlying insight behind them is that pre-trained LLMs (such as Llama3, GPT-3.5 and GPT-4.0, etc.) have amassed a wealth of knowledge that equips them to fulfill various reasoning tasks. However, we observed that merely prompting LLMs to recall experiences related to the target queries may lead to the generation of proxy queries that are as difficult as the target queries themselves, along with corresponding incorrect proxy solutions, misguiding the resolution of the target queries, as exemplified in Fig. 1. See the related works in Appendix A for more details.
The issues mentioned above motivate us to propose a novel automatic reasoning prompt approach, coined LBS3, which is inspired by curriculum learning that mirrors the progressive nature of human learning styles. The idea of curriculum learning (Bengio et al., 2009; Cornacchia and Mossel, 2023) has been widely applied in the field of artificial intelligence, emulating the human learning process from easy to hard tasks (Campos, 2021; Maharana and Bansal, 2022; Huang et al., 2020; Kong et al., 2021; Zhu et al., 2022; Li et al., 2021; Soviany et al., 2022; Xu et al., 2020). Thus, LBS3 critically aims to 1) guide LLMs to generate easy- (or hard-) proxy queries related to the target query and 2) enhance the effectiveness of the solutions for these proxy queries to benefit that for the target query. For the former, diverging from existing approaches (e.g., Self-ICL and Auto-ICL) that generate proxy queries in one pass, we develop a two-stage framework of generation for proxy queries. Specifically, we first prompt LLMs with instructions like "Generate $n_{1}$ different new relevant problems that are easier to solve than the example problem below." to generate simpler proxy queries than the given query, that is, easy-proxy queries. Then, we instruct LLMs to formulate analogical proxy queries for the given query, which are the hard-proxy queries, by using instructions like "Generate $n_{2}$ different new problems that are analogous to the example problem below." Note that $n_{1}$ and $n_{2}$ denote the number of proxy queries generated. For the latter, we initially leverage LLMs to solve each easy-proxy query independently in the ZS-CoT manner, creating corresponding triplet exemplary prompt. Subsequently, we combine these prompts with each hard-proxy query and generate solutions in the FS-CoT fashion. Ultimately, we amalgamate all constructed exemplary prompts with the given query and derive the target solution in the FS-CoT manner. We modularly outline the generic framework of the reasoning process for LBS3 in Fig. 2.
One of the advantages for the proposed approach is that LBS3 explicitly distinguishes between easy- and hard-proxy queries, and ensures that the difficulty of solving proxy queries does not exceed that of the given query. Additionally, in contrast to existing approaches that tackle each proxy query from scratch, we adopt a progressive strategy to harness exemplary prompts derived from easy-proxy queries to guide the generation of solution for hard ones, thereby alleviating the accumulation of errors arises from reasoning ab initio (Yu et al., 2024b). To the best of our knowledge, our work is the pioneering attempt to emulate the idea of curriculum learning, aiming to investigate how LLMs can self-generate few-shot exemplary prompts to facilitate the reasoning process.
@@ 2 Preliminaries
Algorithm 1 The pseudocode of LBS3 approach given one query $[q]$.
1:Input: $[q]$: the target query, $LLM$: large language model, $n_{1}$: the number of easy-proxy queries, $n_{2}$: the number of hard-proxy queries
2:Initial modules: SPG, APG and RAG
3:# Stage 1:
4:prompt_spg=SPG.format($n_{1}$, problem=$[q]$)
[…]
6:exem_sa = {}
7:for $[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]$ in $\{[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]\}_{i=1}^{n_{1}}$ do
8:     prompt_rag-z = RAG-Z.format(problem =$[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]$)
9:     $[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{i}}]$= $LLM($prompt_rag-z))))
10:     exem_sa = exem_sa $\bigcup$ $\{[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}][{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{i}}]\}$
11:# Stage 2:
12:prompt_apg = APG.format($n_{2}$, problem=$[q]$)
[…]
18:# Stage 3:
19:prompt_rag-f = RAG-F.format(exemplars=exem_sa, problem=$[q]$)
20:$[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a}]$ = $LLM($prompt_rag-f))))
21:Output: $[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a}]$
@@ 3.2 Progressive Strategy of Solving Queries
The primary advantage of the above strategy is its effectiveness in enhancing the solutions for hard-proxy queries. To be specific, the easy-to-solve $\{[{\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}q_{i}}]\}_{i=1}^{n_{1}}$ ensures that the corresponding exemplary prompts may be correct with high confidence. Meanwhile, (Chen et al., 2023; Yang et al., 2023; Yasunaga et al., 2023) have shown that using solved proxy queries related or analogous to $[q]$ as exemplary prompts can effectively improve the solution for $[q]$. However, when the difficulty of solving query $[q]$ is high, the generated proxy queries are likely to have comparable challenging, resulting in low-quality exemplary prompts. Therefore, adopting our proposed progressive strategy can alleviate the cumulative errors associated with solving hard-proxy queries from scratch (Yu et al., 2024b), thereby enhancing the quality of their solutions. In addition, we find that using already solved hard-proxy queries as additional exemplary prompts for solving the next hard-proxy query can further strengthen the solution to $[q]$, see Alg. 1 and Section F for more details and empirical justification.
@@
LBS3 94.9 64.2 93.5 86.6 86.0 79.8 92.8 98.0 86.97
4 Experiments
4.1 Experimental Settings
4.2 Results Comparison
Furthermore, the baselines Self-ICL and Auto-ICL uniformly beat ZS-CoT in terms of average accuracy and surpassed FS-CoT in most cases. This result highlights that guiding LLMs to autonomously generate proxy exemplars relevant to a given query can effectively improve their reasoning capabilities. Additionally, the baseline Ana-Pro consistently underperforms other competitors w.r.t. average accuracy, including ZS-CoT, on open-source LLMs, yet consistently outstrips ZS-CoT w.r.t. average accuracy on closed-source LLMs. The said result confirms the high requirements imposed by the Ana-Pro approach on LLMs for following instructions and generating responses.
4.3 Ablation Study
We carefully demonstrate the efficacy and indispensability of the core components in our approach on Qwen1.5-14B-Chat, Llama3-70B-Instruct, and GPT-3.5-turbo over diverse benchmarks. Due to space constraints, we further explore LBS3 with Self-ICL and Auto-ICL and Utility of Progressive Strategy in Appendices E and F, respectively.
4.3.1 Comparison of the number for easy and hard proxy exemplars
We look into the impacts of different hyperparameter combinations ($n_{1}$, $n_{2}$) within the two-stage framework for self-generating proxy queries of LBS3 across various benchmarks, including GSM8K, SQA, CQA, BBH-idfo, BBH-raco, and BBH-ts. For clarity, assume that the number of proxy exemplars $n$ is $3$, with both $n_{1}$ and $n_{2}$ taking values from $\{0,1,2,3\}$. Since $n=n_{1}+n_{2}$, we opt to only consider $n_{1}$, then $n_{2}=n-n_{1}$. When $n_{1}=1$ and $n_{2}=2$, this indicates that LBS3 is tasked with generating one easy-proxy exemplar and two hard-proxy ones. We report the performance of LBS3 corresponding to varying $n_{1}$ on different models and benchmarks in Fig. 3.
@@ 4.3.1 Comparison of the number for easy and hard proxy exemplars
4.3.2 Study for easy- and hard-proxy queries
Based on Section 4.3.1, we further investigate the quality of proxy queries generated by SPG and APG prompt modules and their solution accuracy. To this end, we select the generations of Llama3-70B-Instruct over benchmarks GSM8K, SQA, and BBH-raco. However, determining the quality of generated proxy queries and the correctness of their answers is challenging. On the one hand, there is a lack of effective criteria for dividing the difficulty of generating proxy queries, and on the other hand, there are no plug-and-play standard answers to evaluate the solutions of proxy queries. To overcome the above difficulties, we combine GPT-4o [^3] https://openai.com/api/ and human evaluation to study $50$ randomly sampled target queries in each experiment. Specifically, we first use GPT-4o as a discriminator to score the difficulty of proxy queries generated by SPG and APG in LBS3, with target queries as a reference, following the scoring rules detailed in Table 3 of Appendix D. Then, we invite 10 human participants (all of whom are Ph.D. or Master students) to assess the correctness of the solutions for the generated proxy queries. After completing the above process, we report the average difficulty scores and corresponding solution accuracy of the proxy queries generated by SPG and APG respectively, as shown in Table 2.
Benchmarks $n_{1}$ value num. $n_{1}$/$n_{2}$ SPG-diff APG-diff SPG-acc APG-acc
GSM8K $n_{1}=0$ 0/150 - 0.14 - 92.6
$n_{1}=1$ 50/100 -1.12 0.29 100.0 94.0
$n_{1}=2$ 100/50 -1.47 0.06 100.0 96.0
$n_{1}=3$ 150/0 -1.23 - 100.0 -
SQA $n_{1}=0$ 0/150 - 0.40 - 83.3
$n_{1}=1$ 50/100 -1.60 0.23 92.0 88.0
$n_{1}=2$ 100/50 -1.76 0.16 95.0 90.0
$n_{1}=3$ 150/0 -1.84 - 96.0 -
BBH-raco $n_{1}=0$ 0/150 - 0.03 - 78.7
$n_{1}=1$ 50/100 -1.37 0.02 90.0 80.0
$n_{1}=2$ 100/50 -1.50 0.00 94.0 84.0
$n_{1}=3$ 150/0 -1.14 - 86.9 -
Table 2: Quality study of easy- and hard-proxy queries. Note that, SPG/APG-diff (acc) represents the average difficulty score (solution accuracy(%)) of proxy queries generated by the SPG/APG.
From Table 2, it can be observed that SPG-diff is less than $-1$ and APG-diff is greater than or equal 00 across all benchmarks. Meanwhile, SPG-acc consistently and significantly outperforms APG-acc in terms of accuracy. This indicates that the SPG prompt module can effectively generate proxy queries that are simpler than the target queries (i.e., easy-proxy queries), while the APG prompt module can notably generate proxy queries whose solution difficulty is not lower than that of the target queries (i.e., hard-proxy queries). In other words, compared to the target queries, the easy-proxy queries in LBS3 are indeed simpler, and the solution difficulty of the hard-proxy queries has not decreased and may even have the potential to increase. The aforementioned results confirm the feasibility of LBS3 performing reasoning tasks in a curriculum learning manner. Also, one can see that easy-proxy queries in the LBS3 can effectively augment solution of hard-proxy queries, which in turn improves the reasoning performance for the target queries.